OpenStudy (anonymous):

What is the 8th term of the geometric sequence 4, -20, 100

OpenStudy (anonymous):

\[\LARGE q=\frac{-20}{4}=\frac{100}{-20}=-5\] now you have: \[\LARGE a_8=a_1\cdot q^{8-1}\] \[\LARGE a_8=4\cdot (-5)^{7}\]... pretty much the same as your previous post. ;)